### v1

@@ 5.1 Choosing experimental participants
As a first exercise, we show how observational evidence can inform whom to recruit into an experiment. We consider the setting in Banerjee et al. (2024), who study how the expansion of microfinance affects village social networks in Karnataka, India.
@@
In their observational analysis, the authors assemble panel data from 75 villages in Karnataka, 43 of which were exposed to microfinance. Because program rollout was not randomized across villages, they estimate effects using difference-in-differences (DiD). While DiD is informative, the lack of randomized assignment may bias estimates in the presence of selection and lack of parallel-trends (e.g., Marx et al., 2024; Ghanem et al., 2022). Motivated by these limitations, the authors subsequently conducted an experimental evaluation in one metropolitan area, randomizing microfinance access across 104 urban neighborhoods. We only use observational variation for choosing the experimental design and estimator. We then use experimental variation generated by Banerjee et al. (2024) to validate our procedure.
Using preliminary observational estimates from Banerjee et al. (2024) in Karnataka, we ask: Which area(s) in Karnataka should be prioritized for an experimental evaluation, and how many villages should be enrolled? Concretely, we consider designs that randomize the introduction of microfinance in $n$ villages within one or more selected areas and study how the choice of areas and the allocation of $n$ vary with design constraints. Our objective is to design an experiment that yields an externally valid estimate of the effect of microfinance for rural Karnataka. We focus on the first outcome reported by Banerjee et al. (2024) corresponding to the density of the network, i.e., the percentage of connections a random household in a village has relative to the village size.
In practice, implementation costs often depend on how geographically dispersed the study sites are. Coordinating with microfinance partners, training and supervising field teams, and conducting surveys become more expensive when sites are far apart or span multiple administrative units. We therefore consider operational constraints that limit the dispersion of enrolled villages and number of villages enrolled in the area.
@@
We use observational evidence from Banerjee et al. (2024) to construct area-level difference-in-differences (DiD) estimates for Karnataka. Specifically, we partition the observational sample into four geographically contiguous areas that group nearby villages; each area contains 11–12 villages that were exposed to microfinance during the study period. Figure 4 maps these four areas and reports the corresponding DiD point estimates. Three of the four areas display negative estimated effects, with meaningful variation in magnitudes across areas.
Table 4 summarizes, for each area, the number of treated and untreated villages, the pre-treatment sample variances, and the DiD point estimates together with their variances, computed assuming independence across villages. [^10] Alternative variance estimators (e.g., allowing within-area correlation) can be incorporated without changing the design logic.
These area-specific DiD estimates and their variances serve as our primary observational inputs to the experimental design problem. [^11] We use DiD estimates to mimic the estimator used by the authors. An analogous analysis can be conducted after empirical-Bayes shrinkage of area-level estimates, omitted for brevity.
Table 4: Area-level observational inputs (pre-prior) for Karnataka villages. The outcome is network density (corresponding to the average share of connection of an individual to other individuals in a village). For each area, we report the numbers of treated ($n_{1}$) and untreated ($n_{0}$) villages, the pre-intervention variance of density pooled across arms ($v_{\text{pre}}^{2}$), the area-specific DiD estimate $\hat{\mu}$, and its sampling variance $\widehat{\sigma}^{2}$.
@@
Results
We begin by studying which areas are selected and how sample is allocated when the total experimental sample is large ($n_{1}=52$) and the number of admissible areas may or may not be constrained. Figure 5 visualizes the resulting allocation across Karnataka’s four observational areas for $E\in\{1,2,3,4\}$. With $E=1$, the algorithm concentrates recruitment in the Tumakuru area—the location with the highest uncertainty in the observational estimates. As $E$ relaxes to $2$ and $3$, additional areas enter and the total sample splits across them in roughly—but not exactly—equal shares, reflecting a trade-off between (i) exploiting heterogeneity in the observational variance and (ii) keeping experimental variance low. When $E=4$, all areas are eligible and the allocation smooths further across space. In the rest of our discussion, we discuss properties when not all areas may be selected (E<4E<4). [^12] The solution for $E=4$ always favors minimizing the bias first, since $\beta^{\star}=0$.
Figure 6 tracks how area-level allocations evolve with the total number of villages. When the experiment is small (e.g., $10$ villages) and $E=1$, the algorithm favors an area with relatively low experimental variance—even if its observational estimate is less uncertain—because, at small $n$, experimental noise is first-order. As the total sample increases, the weight shifts toward areas with more uncertain observational estimates. Consistently, the area with the largest experimental variance (Area 2) is typically excluded unless all four areas can be chosen.
Figure 7 reports the optimal shrinkage $\gamma_{j}^{\star}$ by area as total villages vary. The solution is interior for all $E<4$: $\gamma_{j}^{\star}$ rises with sample size, approaching one from below as experimental noise diminishes. For example, at a moderate sample (around 15 villages), $\gamma^{\star}$ is about $0.65$ when $E\in\{2,3\}$; with $E=1$, concentrating the sample within one area lowers experimental variance and therefore yields a larger $\gamma^{\star}$.
Main MSE comparison
@@ Results
Figure 8 plots the ratio $\text{MSE}/\text{MSE}^{\star}$, where $\text{MSE}^{\star}$ is the oracle’s MSE. As the number of treated villages $n_{1}$ increases, all designs improve; because the denominator also falls with $n_{1}$, the ratio need not be monotone. Across $E\in\{1,2,3\}$, the random benchmark remains well above the oracle—by roughly 3–5 percentage points (pp) for $E=1$, about 10 pp for $E=2$, and around 20 pp for $E=3$ at the top of the grid—whereas our design tracks the oracle closely (within a few percentage points) even without knowledge of $b$. The gap is the largest when a larger number of areas $E$ can be included, because our procedure can better emulate the oracle by selecting (and differentially weighting) favorable areas. On the other hand, the random benchmark ignores information about observational and experimental variance heterogeneity. As sample size increases, the gap between the two stabilizes: with large $n_{1}$, optimal shrinkage approaches $\gamma^{\star}\!\approx\!1$ and an even allocation across the selected areas is nearly optimal once the area set is well chosen.
This empirical calibration highlights the benefit of the procedure that is able to track the oracle even in the absence of knowledge of the bias $b$, while significantly outperforming standard alternative benchmarks such as a random allocation strategy.
5.2 Choosing treatment arms for model estimation
Next, we illustrate how our framework selects among treatment arms when the goal is to estimate and deploy a structural model. Combining experimental or pre-program observational data with structural models is increasingly used for program evaluation (Todd and Wolpin, 2006; Attanasio et al., 2012; Meghir et al., 2022). Because large-scale experimentation may be infeasible due to cost or fairness constraints (Muralidharan and Niehaus, 2017), researchers are often interested in designing small-scale experiments whose information, combined with a model, is informative about general-equilibrium (GE) effects.
We consider a researcher evaluating a conditional cash-transfer (CCT) for sending children to school in rural Kenya where such a program is not in place. Due to budget and feasibility constraints, the researcher can only randomize at small scale (partial equilibrium), but ultimately wishes to predict large-scale GE effects. As preliminary observational estimates, we use information from the Mexican PROGRESA experiment.
5.2.1 Model
Individual choice model
Following Bonhomme and Weidner (2022) (and in turn Todd and Wolpin (2006)), let $S\in\{0,1\}$ denote school attendance, $C$ consumption, $Y$ (pre-transfer) household income, $W$ the child’s potential wage, and $t$ the stipend when enrolled. Abstracting from covariates for exposition (we will introduce covariates in the estimation), utility is
with budget C=Y+W(1S)+tSC=Y+W(1-S)+tS. The parametrization $(\xi_{3}-\xi_{1}-\xi_{2})$ is without loss and simplifies expressions below. Enrollment satisfies
$S=1\big\{U(Y+t,1,t,\varepsilon)>U(Y+W,0,0,0)\big\}.$
Letting $Z(Y,W,t)\equiv\xi_{1}W-\xi_{2}Y-\xi_{3}t-\xi_{4},$ we have $P(S=1\mid Y,W,t)=\Phi\!\big(-Z(Y,W,t)\big)$.
General equilibrium effects
We are interested in the effect of a small stipend $t$ to all eligible (poor) households in rural Kenya. GE feedback is allowed through income and wages: for functions $y(t),w(t)$ and mean-zero idiosyncratic income and wage shocks $\varepsilon_{Yi},\varepsilon_{Wi}$, we assume
which decomposes into the direct stipend effect $\phi_{0}\xi_{3}$ and the indirect GE effects via income and wages, $\phi_{0}\xi_{2}y_{0}$ and $\phi_{0}\xi_{1}w_{0}$.
Income and wage effects
We estimate $y_{0}=1.5$ using a local fiscal/income multiplier $M_{\text{tot}}\approx 2.5$ from Egger et al. (2022), which in turn implies $y_{0}=M_{\text{tot}}-1$. [^13] Egger et al. (2022) report a total income/consumption multiplier $M_{\text{tot}}=\partial_{t}\mathbb{E}\big[Y_{\text{pre}}(t)+t\big]|_{t=0}$; therefore the target derivative of pre-transfer income is $y_{0}=\partial_{t}\mathbb{E}\big[Y_{\text{pre}}(t)\big]|_{t=0}=M_{\text{tot}}-1$. We write $w_{0}$ as a function of the parameters: Under simple market clearing for labor supply and demand, we can show [^14] This follows from the following: set $L_{s}(W,t)=L_{d}(W)$, the labor supply equals the labor demand, and differentiate at $t=0$: $\frac{\partial L_{s}}{\partial W}w_{0}+\frac{\partial L_{s}}{\partial t}=\frac{\partial L_{d}}{\partial W}w_{0}$. Suppose we have a constant number of hours worked for working child, denoted as $H$, and any child not at school is working. From the probit index, $\partial_{t}\mathbb{E}[S\,|\,\cdot]=\phi(-Z)\,\xi_{3}$ and $\partial_{W}\mathbb{E}[S\,|\,\cdot]=-\phi(-Z)\,\xi_{1}$ at $t=0$. Hence $\partial L_{s}/\partial t=-\phi_{0}\xi_{3}H$ and $\partial L_{s}/\partial W=\phi_{0}\xi_{1}H$. Dividing the equation by $H$ we obtain the desired result.
$w_{0}=\;\frac{\phi_{0}\,\xi_{3}}{\phi_{0}\,\xi_{1}-d},$ (13)
where $d$ denotes the slope of the demand curve divided by total hours of work. We calibrate $d$ to $0.5\frac{1-S_{0}}{W_{0}}$, where $S_{0}$ and $W_{0}$ are baseline probability to go to school (assuming all children not in school go to work) and baseline wages calibrated to the pre-experimental data from PROGRESA and $0.5$ obtained form the labor demand elasticity in Espey and Thilmany (2000). For expositional convenience, we treat $y_{0},d$ as constant, while $\phi_{0}\xi_{3}$ and $\phi_{0}\xi_{1}$ are parameters of the model and $w_{0}$ is a (non-linear) function of such parameters (it is possible to treat also $y_{0},d$ as additional possibly misspecified parameters, omitted for brevity). [^15] We treat $d$ as a fixed constant given the extensive literature on estimating demand elasiticities, and the fact that $S_{0}$, $W_{0}$ can be calibrated to pre-experimental data (for simplicity here using the PROGRESA pre-experimental data), and $y_{0}$ because computed experimentally in Kenya. When we are concerned with misspecification of those, researchers can augment $\theta$ to also incorporate $y_{0},d$ without having to estimate those estimated in the experiment.
Map to the design parameters
We can now define the parameters of interest as
with
$\tau(\theta)\;=\;\theta_{2}\;+\;y_{0}\,\theta_{1}\;+\;w_{0}(\theta)\,\theta_{3},\qquad w_{0}(\theta)\;=\;\frac{\theta_{2}}{-\theta_{3}-d},$ (15)
where the explicit form of $w_{0}(\theta)$ follows from (13). Given observational estimates $\tilde{\theta}^{\mathrm{obs}}=(\tilde{\theta}^{\mathrm{obs}}_{1},\tilde{\theta}^{\mathrm{obs}}_{2},\tilde{\theta}^{\mathrm{obs}}_{3})$ described below, and following Section 4.1, we use $\omega=\frac{\partial\tau}{\partial\theta}\Big|_{\theta=\tilde{\theta}^{\mathrm{obs}}}$. [^16] Because $w_{0}$ is a function of $\theta$, we have $\frac{\partial\tau}{\partial\theta_{1}}=y_{0},\frac{\partial\tau}{\partial\theta_{2}}=\frac{d}{\theta_{3}+d},\frac{\partial\tau}{\partial\theta_{3}}=-\,\frac{\theta_{2}\,d}{(-\theta_{3}-d)^{2}}.$
5.2.2 Observational study and design of the experiment
Using experimental data from Mexico’s PROGRESA program, we estimate $(\xi_{1},\xi_{2},\xi_{3},\xi_{4})$ via probit with standard controls (age, distance to school, eligibility, year, highest grade attained). Estimation is conducted separately by gender, focusing here on the effects on female students. We pool treated and control observations to estimate the schooling effect $\xi_{4}$, controlling for the individual-level subsidy. In turn, we obtain observational estimates $\big(\tilde{\theta}_{1}^{\mathrm{obs}},\tilde{\theta}_{2}^{\mathrm{obs}},\tilde{\theta}_{3}^{\mathrm{obs}}\big)$ that map to the model parametrization. Standard errors are constructed using the Delta method with clustering at the village level. Point estimates, estimated sensitivity $\omega$, and standard errors are reported below.
Parameter $\tilde{\theta}^{\mathrm{obs}}$ $\omega$ $\sigma$
[…]
We observe heterogeneity in both the sensitivity weights $\omega$ and (more significantly) the precision of the observational estimates. Our goal is to trade off these features, together with prospective experimental variation, to design the experiment in Kenya. Because the setting differs in time and country, PROGRESA-based estimates provide informative but possibly biased baselines.
Candidate small-scale treatment-arms in Kenya (partial equilibrium)
We consider a researcher who can afford only small, partial-equilibrium experiments in Kenya, as they may not be able to randomize treatments across the full population of interest. We examine three possible experiments for a given sample size $n$:
    with precision $v_{1}^{2}/n$. The researcher optimizes over $\gamma_{1}$, the weight used to combine the experimental and observational estimates of $\theta_{1}$, while $\theta_{2}$ and $\theta_{3}$ are calibrated to the PROGRESA study.
    with precision $v_{2}^{2}/n$. The researcher optimizes over $\gamma_{2}$, the weight used to combine the experimental and observational estimates of $\theta_{2}$, while $\theta_{1}$ and $\theta_{3}$ are calibrated to the PROGRESA study.
We calibrate $v_{1}^{2}$ as $\sigma_{1}^{2}\times n^{\mathrm{obs}}$, where $\sigma_{1}^{2}$ is the variance of $\tilde{\theta}_{1}^{\mathrm{obs}}$ and $n^{\mathrm{obs}}$ is the observational-sample size ($n^{\mathrm{obs}}=1089$ for the female sample using data from eligible students in Todd and Wolpin (2006)). We calibrate $v_{2}^{2}$ analogously as $\sigma_{2}^{2}\times n^{\mathrm{obs}}$.
Experiment type Identified parameter Choice variables Experiment standard error Description
5.2.3 Results
Figure 9: Regret–optimal sample allocation across treatment arms when both the unconditional cash transfer (UCT; income shock) and the conditional cash transfer (CCT; stipend) arms are available. The optimal solution is interior: the vast majority of participants (over 90%) are assigned to CCT, with a small but nonzero fraction assigned to UCT to hedge misspecification. This pattern reflects the higher payoff to learning about CCT in this application relative to UCT.
We now summarize results as we vary both the set of experiments the researcher can run and the total sample size $n$. Specifically, Figure 9 reports the sample allocation implied by the regret-optimal design when both treatment arms are available. The optimal allocation is interior: most participants (over roughly $90\%$) are assigned to the CCT arm (education), with a small but nonzero fraction assigned to the UCT arm (income shock). Two forces rationalize this pattern. First, the CCT parameter is more misspecification-sensitive ($\omega_{2}\approx 1.98$ vs. $\omega_{1}\approx 1.5$), so generating experimental evidence on CCT has a larger payoff in bias reduction. Second, the income effect $\theta_{1}$’s variance is about an order of magnitude smaller than the stipend effect $\theta_{2}$’s variance in our PROGRESA calibration, lowering the marginal returns to learn about UCT relative to CCT. Hence, while the designer keeps some allocation on UCT, most of the sample is placed on CCT; our framework makes this bias–variance trade-off explicit.
Alongside the sample allocation, our method optimizes the shrinkage weights on the parameters of interest. Figure 10 (left panel) shows that when both arms can be run, we place essentially full weight on the experimental estimator for UCT ($\gamma^{\star}_{\text{UCT}}\approx 1$ across $n$), while for CCT we gradually shift toward the experimental estimator as precision improves: $\gamma^{\star}_{\text{CCT}}$ rises from about $0.5$ at small $n$ to approximately $0.9$ by $n=1000$. Intuitively, higher $n$ lowers experimental variance, making it optimal to rely more heavily on experimental evidence for the CCT parameter.
It is also informative to compare optimal choices when the researcher is restricted to one arm. Figure 10 (right panel) displays the selected arm as a function of $n$ (together with the corresponding $\gamma_{j}^{\star}$). We encode “not chosen” as $\gamma_{j}^{\star}=0$. For small samples ($n\lesssim 500$), the optimal choice is UCT: the experimental CCT estimator would be too noisy, so the resulting variance under CCT dominates, while the bias advantage of CCT over UCT is limited because $\omega_{1}$ and $\omega_{2}$ are of similar magnitude. Once $n$ crosses a critical threshold (around $n\approx 500$), CCT becomes dominant: by combining observational and experimental estimates for CCT, we can achieve comparable variance, and the worst-case bias is lower than under UCT.
Finally, Figure 11 reports worst-case regret for each design choice. The left panel shows that the regret of our method (green line, running both arms with an interior allocation) declines toward one as $n$ increases, reflecting near-oracle performance under our normalization. In contrast, single-arm designs exhibit nonvanishing regret because bias remains first order even as variance shrinks. The right panel focuses on the single-arm problem: both the oracle and the researcher can only choose a single arm. In this case, UCT is preferable only at small $n$; beyond the same threshold ($\approx 500$), CCT yields uniformly lower regret as variance becomes second order relative to bias. The regret vanishes relative to the oracle that can only choose a single arm.
These results illustrate how our approach systematically balances precision and misspecification risk to balance experimental variation with structural models.
@@ 6 Implications for practice
This paper studies experimental design in the presence of complementary observational evidence. Practitioners often have access to observational inputs—e.g., estimates from models trained on existing data (which may be misspecified), results from prior experiments that may not transport to the target context, or estimates from observational designs that may suffer from confounding. Our goal is to combine such evidence with new experimental data, which deliver unbiased estimators for specific parameters of interest. We ask which experiment to run, and how to run it, under budget constraints that limit the number of arms and/or impose per-unit (variable) costs on effective sample size.
A key challenge is that the bias of the observational estimators is unknown in practice. We adopt a minimax proportional regret criterion that compares the mean-squared error (MSE) of a candidate design to that of an oracle that knows the worst-case bias. This reveals a fundamental trade-off between precision and robustness. The optimal design balances the design’s variance normalized by the smallest achievable variance (variance gap) and its worst-case bias normalized by the smallest attainable bias (bias gap). We propose a procedure that jointly determines: (i) how to combine observational and experimental evidence; (ii) how to allocate precision across experiments given budget constraints; and (iii) which treatment arm and/or sub-population to include in the experiment with fixed experimental costs.
@@
  • Define the estimand(s) of interest. Specify $\tau(\theta)$ for a known mapping $\tau$ and unknown parameters $\theta\in\mathbb{R}^{p}$. For example, $\tau(\theta)$ may represent a counterfactual, a general-equilibrium effect, or an average impact across locations. Adopt a parametrization in which some (but not necessarily all) components of $\theta$ can be learned experimentally; this clarifies what the experiment can identify.
  • Assemble informative observational evidence. Collect observational estimates $\tilde{\theta}^{\rm obs}$ and their covariance $\tilde{\Sigma}^{\rm obs}$. These serve as informative (but potentially biased) baselines. Such evidence may come from observational designs, structural estimates with pre-experimental data, or prior experiments conducted in different contexts or periods.
@@
  • Specify feasibility constraints and calibrate experimental variance. Enumerate the admissible design set $\mathcal{S}$. That is, $\mathcal{S}$ is the set containing combinations of parameters that can be learned jointly via experiments (which arms, sites, or mechanisms can be randomized). Define, for each combination of parameters $\mathcal{E}\in\mathcal{S}$, the corresponding precision set $\mathcal{G}(\mathcal{E})$ as the feasible set of experimental variances for a given experiment $\mathcal{E}$. This requires specifying how the sample size $n_{j}$ allocated to each arm maps to the variance of each experimental estimate. Calibrate per-unit experimental variances using pilot studies or historical data.
  • Run the method and decide. For each feasible set of experiments $\mathcal{E}\in\mathcal{S}$ and precisions $\Sigma\in\mathcal{G}(\mathcal{E})$, we first optimize over the shrinkage vector $\gamma$ (how to weight observational vs. experimental estimates). We then optimize over the precision/allocation $\Sigma$ (e.g., the sample size $n_{j}$ assigned to each arm). Finally, we optimize over the experiment choice $\mathcal{E}\in\mathcal{S}$. In practice, this involves enumerating feasible $\mathcal{E}$ and simple numerical optimization for $(\gamma,\Sigma)$. The output is a pre-analysis plan with the selected arm(s), sample sizes, and pre-specified combination rule ($\gamma$).
  • Reporting and diagnostics. In a pre-analysis plan, we recommend reporting (i) the estimands of interest; (ii) the chosen arm(s) and the final sample size allocation, (iii) the shrinkage weights $\gamma^{\star}$ by parameter between the observational and experimental study, and (iv) the two normalized components $\alpha/\alpha^{\star}$ (variance-gap) and $\beta/\beta^{\star}$ (bias-gap), that can help assess the quality of the allocation. Before committing to a single experiment choice, researchers may also want to explore sensitivity to a range of values for the experimental variances.
@@
The applicability of our framework spans a wide range of settings in economics and beyond. Examples include estimating general equilibrium effects or structural models (Todd and Wolpin, 2006; Attanasio et al., 2012; Meghir et al., 2022; Kreindler et al., 2023; de Albuquerque et al., 2025); choosing among alternative treatment arms in factorial designs (Muralidharan et al., 2020; Bandiera et al., 2025); and deciding where to run the next experiment for external validity (Gechter et al., 2024; Olea et al., 2024). In industrial organization, applications include choices between demand-side and supply-side interventions (Bergquist and Dinerstein, 2020), the effects of information acquisition in markets (Allende et al., 2019; Larroucau et al., 2024), and decisions about which additional data source to acquire to improve statistical analysis. Beyond economics, medical applications include allocating sample size across subgroups and allocating doses across treatment arms (e.g., Porter et al., 2024; Morita et al., 2017; Manski, 2025). In all cases, the same recipe identifies which parameters to target, how much to learn about each, and how to blend experimental and observational evidence.
@@
Several open questions remain for future research. These include settings where researchers have well-specified priors about bias (e.g., from meta-analysis), sequential or adaptive experimental choices, and environments where the exogenous source of variation yields only partial identification of some parameters of interest.
Appendix A Additional figures
Appendix B Proofs of main results

### v2

@@ 6.2 Measuring the performance of our method for site selection
Next, we show how observational evidence can inform whom to recruit into an experiment. We consider the setting in Banerjee et al. (2024), who study how the expansion of microfinance affects village social networks in Karnataka, India. We focus on the first outcome reported by Banerjee et al. (2024) corresponding to the density of the network, i.e., the percentage of connections a random household in a village has relative to the village size.
Background
@@
In their observational analysis, the authors assemble panel data from 75 villages in Karnataka, 43 of which were exposed to microfinance. Because program rollout was not randomized across villages, they estimate effects using Difference-in-Differences (DiD). While DiD is informative, the lack of randomized assignment may bias estimates in the presence of selection (e.g., Ghanem et al., 2022). Motivated by these limitations, the authors subsequently conducted an experimental evaluation in one metropolitan area, randomizing microfinance access across 104 urban neighborhoods. We only use observational variation for choosing the experimental design and estimator. We then use experimental variation generated by Banerjee et al. (2024) to validate our procedure.
Research question
In practice, implementation costs often depend on how geographically dispersed the study sites are due to coordination and survey costs. Using preliminary observational estimates from Banerjee et al. (2024) in Karnataka, we ask: Which area(s) in Karnataka should be prioritized for an experimental evaluation, and how many villages should be enrolled?
@@
We partition the observational sample into four geographically contiguous areas that group nearby villages; each area contains 11–12 villages that were exposed to microfinance during the study period. Figure 7 maps these four areas and reports the corresponding DiD point estimates. Three of the four areas display negative estimated effects, with meaningful variation in magnitudes across areas. Table 2 summarizes, for each area, samples sizes and variances, clustered at the village level. These area-specific DiD estimates and their variances serve as our primary observational inputs for experimental design. [^12] We use DiD estimates to mimic the estimator used by the authors. An analogous analysis can be conducted after empirical-Bayes shrinkage of area-level estimates, omitted for brevity.
Table 2: Area-level observational inputs for Karnataka villages. The outcome is network density (corresponding to the average share of connection of an individual to other individuals in a village). For each area, we report the numbers of treated ($n_{1}$) and untreated ($n_{0}$) villages, the pre-intervention variance of density pooled across arms ($v_{\text{pre}}^{2}$), the area-specific DiD estimate $\hat{\mu}$, and its sampling variance $\widehat{\sigma}^{2}$.
@@
Experimental design
We begin by studying which areas are selected and how sample is allocated when the total experimental sample is large ($n_{1}=52$) and the number of admissible areas may or may not be constrained. Figure 8 visualizes the resulting allocation across Karnataka’s four observational areas for $E\in\{1,2,3,4\}$. With $E=1$, the algorithm concentrates recruitment in the Tumakuru area—the location with the highest uncertainty in the observational estimates. As $E$ relaxes to $2$ and $3$, the total sample splits across them in roughly (but not exactly) equal shares, reflecting a trade-off between (i) exploiting heterogeneity in the observational variance and (ii) keeping experimental variance low. When $E=4$, all areas are eligible and the allocation smooths further across space. In the rest of our discussion, we discuss properties when not all areas may be selected ($E<4$).
Figure 9 tracks how area-level allocations evolve with the total number of villages. When the experiment is small (e.g., $10$ villages) and $E=1$, the algorithm favors an area with relatively low experimental variance, even if its observational estimate is less uncertain, because, at small $n$, experimental noise is first-order. As the total sample increases, the weight shifts toward areas with more uncertain observational estimates. Consistently, the area with the largest experimental variance (Area 2) is typically excluded.
Appendix Figure 13 reports the optimal shrinkage $\gamma_{j}^{\star}$ by area as total villages vary. The solution is interior for all $E<4$: $\gamma_{j}^{\star}$ rises with sample size, approaching one from below as experimental noise diminishes. For example, at a moderate sample (around 15 villages), $\gamma^{\star}$ is about $0.65$ when $E\in\{2,3\}$; with $E=1$, concentrating the sample within one area lowers experimental variance and therefore yields a larger $\gamma^{\star}$.
Main MSE comparisons
@@
Figure 10 plots the ratio $\text{MSE}/\text{MSE}^{\star}$, where $\text{MSE}^{\star}$ is the oracle’s MSE. As the number of treated villages $n_{1}$ increases, all designs improve; because the denominator also falls with $n_{1}$, the ratio need not be monotone. Across $E\in\{1,2,3\}$, the random benchmark remains well above the oracle—by roughly 3–5 percentage points (pp) for $E=1$, about 10 pp for $E=2$, and around 20 pp for E=3E=3—whereas our design tracks the oracle closely (within a few percentage points) even without knowledge of $b$. The gap is the largest when a larger number of areas $E$ can be included, because our procedure can better emulate the oracle by selecting (and differentially weighting) favorable areas.
This result illustrates the merit of the method that improves MSE up to $20$ percentage points compared to a standard benchmark and leads to a near-oracle solution.
@@ 7 Implications for practice
This paper studies experimental design in the presence of observational evidence. A key challenge is that the bias of the observational estimators is unknown in practice. We adopt a minimax proportional regret criterion that compares the mean-squared error (MSE) of a candidate design to that of an oracle that knows the worst-case bias. This reveals a fundamental trade-off between precision and robustness. The optimal design balances the design’s variance normalized by the smallest achievable variance (variance gap) and its worst-case bias normalized by the smallest attainable bias (bias gap). We propose a procedure that jointly determines: (i) how to combine observational and experimental evidence; (ii) how to allocate precision across experiments given budget constraints; and (iii) which treatment arm and/or sub-population to include in the experiment with fixed experimental costs.
@@
  • Define the estimand(s) of interest. Specify $\tau(\theta)$ for a known mapping $\tau$ and unknown parameters $\theta\in\mathbb{R}^{p}$. For example, $\tau(\theta)$ may represent a counterfactual, a general-equilibrium effect, or an average impact across locations. Adopt a parametrization in which some (but not necessarily all) components of $\theta$ can be learned experimentally; this clarifies what the experiment can identify. When such parametrization is not readily available, Section 4 shows how our procedure applies to $\tau$ identified through arbitrary moment restrictions.
  • Assemble informative observational evidence. Collect observational estimates $\tilde{\theta}^{\rm obs}$ and their covariance $\tilde{\Sigma}^{\rm obs}$. These serve as informative (but potentially biased) baselines. Such evidence may come from observational designs, structural estimates with pre-experimental data, or prior experiments conducted in different contexts.
@@
  • Specify feasibility constraints and calibrate experimental variance. Enumerate the admissible design set, and the budget on sample size allocation. Calibrate per-unit experimental variances using pilot studies or historical data to form the set of feasible designs $\mathcal{D}$.
  • Run the method and decide. Optimize over the design and corresponding estimators. The output is a pre-analysis plan with the selected arm(s), sample sizes, and pre-specified combination rule ($\gamma$).
@@
The applicability of our framework spans a wide range of settings in economics and beyond. Examples include estimating general equilibrium effects or structural models (Todd and Wolpin, 2006; Attanasio et al., 2012; Meghir et al., 2022; Kreindler et al., 2023; de Albuquerque et al., 2025); choosing among alternative treatment arms in factorial designs (Muralidharan et al., 2020; Bandiera et al., 2025); and deciding where to run the next experiment for external validity (Gechter et al., 2024; Olea et al., 2024). In industrial organization, applications include choices between demand-side and supply-side interventions (Bergquist and Dinerstein, 2020), the effects of information acquisition in markets (Allende et al., 2019; Larroucau et al., 2024), and decisions about which additional data source to acquire to improve statistical analysis (Allcott et al., 2025). Beyond economics, medical applications include allocating sample size across subgroups and allocating doses across treatment arms (e.g., Porter et al., 2024; Morita et al., 2017; Manski, 2025).
@@
Several open questions remain for future research. These include settings where researchers have well-specified priors about bias such as from Bayesian models (Gechter et al., 2024), or other modeling choices (Olea et al., 2024). This may be potentially incorporated through additional moment conditions as in Section 4. Additional future extensions include sequential or adaptive experimental choices or settings where the target estimand is only be partially identified (e.g. Cesa-Bianchi et al., 2025).
Appendix A Proofs of main results
A.1 Proof of Theorem 1 and Theorem 2
Theorem 1 is a special case of Theorem 2 proved below.
Step 1 (worst-case MSE). By Hölder’s inequality,
$\sup_{b\in\mathcal{B}(B)}\big(\omega^{\top}\Gamma_{\Lambda}(W)\,b\big)^{2}\leq\Big(B\,\big\|\big(\Gamma_{\Lambda}(W)^{\top}\omega\big)_{\mathcal{I}^{c}}\big\|_{1}\Big)^{2}=B^{2}\,\beta(W).$
In addition, this inequality is attained by $b_{j}=B\,\mathrm{sign}\!\big((\Gamma_{\Lambda}(W)^{\top}\omega)_{j}\big)$ for $j\in\mathcal{I}^{c}$ (and $b_{j}=0$ for $j\in\mathcal{I}$). Hence $\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(W,\Sigma)=\alpha(W,\Sigma)+B^{2}\beta(W).$
Step 2 (oracle envelope and basic properties). Let $\delta(t):=\inf_{W,\Sigma}\{\alpha(W,\Sigma)+t\beta(W)\}$. Recall that $\beta(W)$ and $\alpha(W,\Sigma)$ are bounded from above and $\alpha(W,\Sigma)>0$ by Assumption 3. Being the pointwise infimum of affine functions of $t$, $\delta$ is concave, nondecreasing, and finite for $t\geq 0$. Moreover,
$\delta(0)=\alpha^{\star},\qquad\lim_{t\to\infty}\frac{\delta(t)}{t}=\beta^{\star}.$
Step 3 (quasi-convexity and boundary maximization). Fix $(\alpha,\beta)$. Consider
$\phi(t):=\frac{\alpha+t\beta}{\delta(t)},\qquad t\geq 0.$
Because the numerator is affine and the denominator is positive and concave in $t$, $\phi$ is quasi-convex on $0,\infty)$. Hence, on any compact interval $[0,T]$, $\max_{t\in[0,T]}\phi(t)$ is attained at the boundary $\{0,T\}$. Letting $T\to\infty$ and using the limits in Step 2,
$\sup_{t\geq 0}\phi(t)=\max\!\left\{\frac{\alpha}{\delta(0)},\ \lim_{t\to\infty}\frac{\alpha+t\beta}{\delta(t)}\right\}=\max\!\left\{\frac{\alpha}{\alpha^{\star}},\ \frac{\beta}{\beta^{\star}}\right\},$
with the stated conventions $0/0=0$.
Appendix B Proofs of the extensions